OpenStudy (anonymous):

Simplify the expression (6-4i)(-1+6i)

OpenStudy (jdoe0001):

FOIL it, keeping in mind that \(\bf i^2\implies \sqrt{-1}\cdot \sqrt{-1}\implies (\sqrt{-1})^2\implies \sqrt{(-1)^2}\implies {\color{blue}{ -1}}\)

OpenStudy (anonymous):

18 + 40i